OpenStudy (anonymous):

Find the solution sets for Ix^2-x+9I<3

OpenStudy (anonymous):

\[|x^2-x+9|<3\] \[-3<x^2-x+9<3\] so you have two inequalities to solve: \[x^2+x-9<3\] and \[-3<x^2-x+9\] and you have to take the intersection

OpenStudy (anonymous):

oh so the 3 changes sign not the other side

OpenStudy (anonymous):

first one: \[x^2-x+9<3\] \[x^2-x+6<0\] \[(x-3)(x+2)<0\]

OpenStudy (anonymous):

yes if \[|blah|<p\] then \[-p<blah<p\]

OpenStudy (anonymous):

oh ok that's where I got stuck I was doing IbahI<3 then -blah>3 or blah<3

OpenStudy (anonymous):

the first one is zero if x =-2 or x = 3 and it is negative between them so the solution is \[(-2,3)\]

OpenStudy (anonymous):

we still have more work to do. (-2,3) is just the solution to the first inequality. you have to solve them separately

OpenStudy (anonymous):

right the other is (3,4)

OpenStudy (anonymous):

I just did it now.

OpenStudy (anonymous):

the second inequality is really?

OpenStudy (anonymous):

How do I write that as an answer?

OpenStudy (anonymous):

that sucks because that means there is no solution

OpenStudy (anonymous):

let me try it hold on

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

sorry 3, -4

OpenStudy (anonymous):

oooh. let me try it anyway

OpenStudy (anonymous):

so its (-4,-4)u(2,3)?

OpenStudy (anonymous):

\[-3<x^2-x+9\] \[0<x^2-x+12\] \[0<(x-4)(x+3)\]

OpenStudy (anonymous):

aaah i see your confusion

OpenStudy (anonymous):

oh woops.

OpenStudy (anonymous):

this inequality says that that product is positive

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

so answer is not (-3,4) but rather \[(-\infty,-3) \cup (4,\infty)\]

OpenStudy (anonymous):

oh ok.

OpenStudy (anonymous):

is that the whole answer?

OpenStudy (anonymous):

so now we have to take the intersection of those two sets.

OpenStudy (anonymous):

which looks empty to me. hold on let me try it again real quick. it is possible there is no solution

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

k we started with \[x^2-x+9\] yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the minimum value of this thing is where \[x=-\frac{b}{2a}\] which in this case is \[\frac{1}{2}\]

OpenStudy (anonymous):

woah woah woah.

OpenStudy (anonymous):

what's that?

OpenStudy (anonymous):

plug in x = 1/2 get y = 35/4=8.75

OpenStudy (anonymous):

which is bigger than 3!

OpenStudy (anonymous):

did i lose you there?

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

ok let me ask you this. do you know what the graph of \[y=x^2-x+9\] look like?

OpenStudy (anonymous):

let me look up the answer in the answer key real quick

OpenStudy (anonymous):

yes, a parabola

OpenStudy (anonymous):

the answer is no solution. yes a parabola that faces upward yes?

OpenStudy (anonymous):

because it faces upward, it has a minimum value, where the parabola "sits" that is called the vertex

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

the vertex of this parabola is \[(\frac{1}{2},\frac{35}{4})\]

OpenStudy (anonymous):

35/4=8.5

OpenStudy (anonymous):

and of course 8.5 > 3

OpenStudy (anonymous):

(a) (−3,−2) ∪ (3,4)

OpenStudy (anonymous):

that's the answer...

OpenStudy (anonymous):

so there is no way that this thing can be less than 3

OpenStudy (anonymous):

no chance

OpenStudy (anonymous):

here is a picture of it. you will see that it is never less than 3 http://www.wolframalpha.com/input/?i=y%3Dx^2-x%2B9

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

the answer is "no solution"

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

we found that in any case. the intersection of the 3 intervals we found is empty

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

one was (-2,3) and the other two were \[(-\infty,-3)\] and \[(4,\infty)\]

OpenStudy (anonymous):

that intersection is empty

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks.